Given the sequence 1331, 1000, 729, x, y, z, 125, . . . , what is the sum of x, y, z? Please show me the steps.
Answer by richard1234(7193)

1331 = 11^3
1000 = 10^3
729 = 9^3
.
.
.
125 = 5^3

Therefore, x,y,z = 8^3, 7^3, 6^3 or 512, 343, 216. Their sum is 512+343+216 = 1071.